a shopkeeper has three varieties of oranges which he sells at the rate 6 rupees for dozen,5 rupees for dozen and 3 rupees for dozen.what is his average rate
1
Expert's answer
2012-03-06T12:00:31-0500
The averge rate is P = (6 + 5 + 3)/3 = 14/3 rupees for dozen.